OpenStudy (anonymous):

Find all zeros by factoring each function. f(x)= x^4 - x^2 - 30

OpenStudy (cwrw238):

= (x^2 + a)(x^2 + b) where a*b = -30 and a + b = -1 can you figure out the values of a and b?